# Runtime Stealthy Perception Attacks against DNN-based Adaptive Cruise Control Systems

### v3

@@ III-A Attack Model
A possible way to achieve this is to implant malware by compromising the over-the-air (OTA) update mechanisms [4, 87, 64, 35] or gaining one-time remote access to the ADAS software through scanning the network, accessing stolen credentials and exploiting the vulnerabilities in SSH protocol [17], browsers, access control [68], wireless communications [68, 57, 38], third-party components connected to in-vehicular network [48], or some remote service/backdoor offered by the manufacturer (e.g., Comma Connect for OpenPilot [3] or Bluelink for Hyundai). For example, a publicly-available tool developed for OpenPilot enables an attacker on the same network as a target device to install a malicious code [6]. With such remote access, the attacker can also change the OTA settings (e.g., remote URL) to prevent potential patches from being effective. This assumption about the attack surface for deploying malware is also supported by previous works [35, 34], and could have a large impact as it can be generalized to any vehicle with similar OTA and DNN mechanisms and target a large fleet of vehicles at the same time.
TABLE II: Threat models: attacker strength, capability, and impact.
Threat Model Attacker Strength Access to ADAS Software Vehicular Networks Impact Examples
Read Write
Malware Strong1 $\checkmark$ $\checkmark$ $\checkmark$
@@
[35, 34]
Wireless Medium2 $\checkmark$ $\checkmark$
@@
[50][72][44]
[68][63]
Physical Weak3 $\checkmark$
@@
[41][25]
[56][62][59]
@@
    Only perception attacks possible
Another way to compromise live camera data is to connect to a wireless communication device, either a third-party component or one implemented by an attacker, connected to the vehicular network, such as ROS communication channels [50], CAN Bus [68, 13, 63, 87] or Ethernet channel [44, 72]), to read and send image data at runtime.
Further, physical attack methods are also viable, such as by displaying the patch on a monitor attached on the rear side of a leading adversarial vehicle [41, 25] or projecting the patch into the rear of the lead vehicle using a projector [56, 62, 59].
Table II summarizes various methods for runtime reading and modifying of live camera frames, given different attacker strengths and capabilities. In this paper, we mainly focus on the runtime and optimized modification of live camera frames to enhance the attack success rate and stealthiness, regardless of the exact threat model and how the attacker obtained access. In our experiments, we implemented the attack through malicious OTA update to OpenPilot as described in Section III-C.
Attacker Constraints. We restrict the scope of attack capabilities in the paper to reading and modifying live camera data, ensuring uniformity across all presented threat models. Although some attack models could potentially enable more aggressive attacks, such as directly altering ACC controller outputs ( 3 in Fig. 1) via malware or wireless method or changing the DNN output ( 2) with malware method, these may be more readily identified by safety mechanisms, such as AEB, or by a human driver, as discussed in Section V-D.
@@ III-B Attack Challenges
C1. Optimal timing of attacks at runtime to cause safety hazards. Prior attacks on ADAS that rely on random strategies to determine the attack timing (start time and duration) have proven ineffective in achieving a high attack success rate [97, 73, 43] as they waste computational resources by trying random attack parameters that lead to no safety hazards. For instance, initiating an attack on an Ego vehicle to induce sudden acceleration does not cause safety hazards when no lead vehicle is detected. Recent works have focused on using machine learning to explore the fault/attack parameter space [66] and improve the attack effectivenes [43, 44], but they still require substantial amounts of data from random attack experiments for model training. Finding the optimal triggering time and duration is crucial for effective attacks, yet challenging due to the vast parameter space that needs exploration.
C2. Generating attack value at runtime to adapt to dynamic changes in the driving environment. Attacking DNN-based ACC systems on a moving vehicle faces challenges due to continuous variations in the driving environment, such as object position and size captured by the Ego vehicle’s camera. Existing attack algorithms [36, 39] are inadequate as they plan perturbations offline, assuming fixed sizes and locations for attack vectors. A new algorithm is needed to dynamically adapt the attack vector’s value (e.g., position, dimension, and amount of perturbation) to match the lead vehicle’s dynamics. These changes disrupt the original attack vector generation process, requiring a unique approach to address inconsistencies and non-differentiability in the objective function. In addition, the attack value should be designed in a stealthy way to avoid detection by the human driver or safety mechanisms.
C3. Incorporating real-time constraints into the attack optimization process. Previous attacks on DNN models assume predetermined target images [36] or a known set [76, 39] with unlimited computation resources, allowing iterative optimization until an optimal attack vector is generated. However, attacking ACC systems in real-time presents challenges as the camera continuously provides frames without prior knowledge. An attack vector must be generated in real-time before the next frame or control action execution. The real-time control cycle and camera update frequency limit the speed of generating the adversarial attack vector and the frequency of assessing the perturbation’s impact on DNN predictions. These tight constraints in typical ACC systems (e.g., frame rate of 20Hz and control cycle of 10 milliseconds in [29]) significantly impact the effectiveness of optimization-based attack strategies.
III-C Attack Design
@@
To tackle challenge C1, rather than randomly or exhaustively exploring the attack parameter space, we concentrate on systematically characterizing specific values within the parameter space (e.g., attack start time and duration). This targeted approach aims to identify optimal system contexts (or critical times) for activating the attacks to not waste time and resources on non-hazardous attack scenarios (Section III-C1).
@@
To address the challenges C2-C3, we design a novel optimization approach and an adaptive algorithm to dynamically determine optimal pixel values for an adversarial patch at runtime, aiming to maximize the error in DNN-based lead vehicle predictions and to accommodate dynamic changes in the driving environment. To optimize attack effectiveness and computational efficiency, our method focuses on the small area of the bounding box around the target vehicle and employs a primary attribution algorithm [83] to identify and manipulate the most crucial pixels (Section III-C2). Additionally, our adaptive algorithm retains optimization results of perturbation size, position, and value from the previous perception cycle instead of restarting the optimization process (contrary to other iterative optimization methods [31]) to satisfy the tight real-time constraints of the perception system (50ms).
Figure 2: Attack design: Offline preparation, Runtime execution.
@@
The attacker initiates the attack process by conducting an offline analysis on the target DNN-based ACC system. The analysis includes examining operational data and open-source code to identify the target software files and functions and/or the DNN input and output fields to be monitored and infected. Then the attack steps outlined in the orange box in Fig. 2, are implemented and executed either remotely or locally through unauthorized access to the target ACC under one of the threat models described in Table II. In Fig 2, we present an attack implementation example via malware method that will replace the target functions or system libraries. The attacker installs the malware on the target ACC system through one-time access to the victim vehicle control system, achieved by exploiting remote access vulnerabilities or compromising OTA updates (Section III-A). Appendix B provides an illustrative example of how we install the malware on an OpenPilot system. At runtime, the malicious code intercepts sensor data before reaching the DNN model and infers the current system context for activating attacks. When the detected system context aligns with a critical system context (Section III-C1), an optimized adversarial patch is generated (Section III-C2) and added to the image data and sent to the DNN model.
III-C1 Context-Aware Attack Activation
@@
To find the most critical times for activation of the attack (C1), we adopt a control-theoretic hazard analysis method [52] to identify the most critical system contexts under which specific control actions are unsafe and, if issued by the ACC control software, could lead to hazards. This approach mainly relies on domain knowledge about system safety requirements and, contrary to an ML-based approach, does not require large amounts of training data or computation resources.
@@
In our hazard analysis, we define the accident as the adverse event of forward collision (measured by a zero or negative relative distance between the Ego vehicle and the lead vehicle or a front object). This can happen as a result of the Ego vehicle transitioning into a hazardous state that violates the safe following distance with the lead vehicle.
@@
To determine the critical system contexts, we assess all the combinations of system states and ACC control actions (e.g., acceleration, deceleration) to identify the specific combinations that are most likely to lead to hazards. Table III shows part of the safety context table for ACC with a focus on the acceleration commands. Here, the critical system context is described as when the Headway Time (HWT, the time the Ego vehicle takes to drive the relative distance ($RD$) from the lead vehicle with the current speed) is less than a safety limit HWTsafe (e.g., 2-3s), and the Ego vehicle is faster than the lead vehicle ($V_{Ego}>V_{Lead}$). Under such system context, an acceleration command induced by the camera perception attacks or other reasons will most likely lead to a forward collision hazard. This high-level specification of critical system context can be done by an attacker based on the knowledge of the typical functionality of an ACC system and be applied to any ACC system with the same functional specification.
TABLE III: Partial safety context table for an ACC system.
@@ 4 RS>0 No
Figure 3: Optimization-based adversarial patch generation.
III-C2 Adaptive Adversarial Patch Generation
@@
min $\sum_{d\in{RD}_{t}}{-\triangledown g(d,\theta)}+\lambda||\Delta_{t}||_{p}$ (1)
s.t. $Patch_{t}=\Delta_{t}*M_{t}$ (2)
$Patch_{t}\in[\mu-\sigma,\mu+\sigma]$ (3)
$Area(Patch_{t})\subset BBox(LV)_{t}$ (4)
$X^{adv}_{t}=X_{t}+Patch_{t}$ (5)
$[{RD},RS]_{t}=LVD_{\theta}(X^{adv}_{t-1})$ (6)
$u_{t}=ACC(s_{t},[{RD},{RS}]_{t})$ (7)
$s_{t+1}=CarModel(s_{t},u_{t})$ (8)
@@
In Eq. 1, $g(d)$ is an approximate polynomial function of $d$ that fits the trend of the trajectory of the relative distance ${RD}_{t}$, predicted by the DNN model with weight parameters $\theta$. "-" is a negative sign that converts our goal of maximizing the relative distance to minimizing the proposed objective function. For example, when the gradient of the relative distance trajectory, $g(d)$, is negative, minimizing "$-\triangledown g(d)$" will slow down the decrease of $g(d)$ and is equivalent to maximizing the relative distance. We adopt the gradient of $g(d)$ in the objective function instead of using ${RD}_{t}$ itself in order to avoid sharp changes in the predicted relative distance value, which might be easily detected by some anomaly detection mechanisms.
@@
In Eq. 5, the perturbation is added to the original image input $X_{t}\in\mathbb{R}^{H\times W\times C}$ in the form of an adversarial patch $Patch_{t}\in\mathbb{R}^{H\times W\times C}$, represented as a matrix of pixels with height H, width W, and C color channels. $\lambda$ is the weight parameter of the p-norm regularization term, designed to minimize the perturbation value of the patch for stealthiness. We limit the perturbation value within the Kalman filter noise parameters ($\mu$,$\sigma$) (Eq. 3), which ensures the perturbation is not corrected by the sensor fusion. We also constrain the adversarial patch inside the BBox of the lead vehicle (Eq. 4) to enhance attack effectiveness, minimize the perturbation area for stealthiness, and reduce computational cost.
@@
Figure 4: Examples of the shift and adjustment process in the patch generation due to the changes in the lead vehicle’s position and size from the driver’s view. Inset figures are the zoomed-in views of the front vehicle with an adversarial patch added around the license plate area.
Figure 5: ACC under attack.
@@
Primary Attribution Detection and Patch Update. As mentioned in Section III-B, a major challenge (C3) in the design of runtime attacks is the changes in the size and location of the lead vehicle in the perceived image frames. To address this challenge, the attacker needs to update the generated patch dynamically according to the approximate DNN outputs. In this paper, we adopt an object detection method [71] to detect and track the real-time location and size of the lead vehicle and concentrate the attack perturbation within the detected BBox of the lead vehicle. In production ACC with object detection features [21], given proper access, the attack can skip this step and use the stock prediction results.
@@
After getting the BBox, we utilize a primary attribution algorithm [83] to quantify the relationship between input features and output predictions. Through this exploration, we try to identify the important pixels within the BBox of the lead vehicle that contribute the most to the predictions of $RD_{t}$. The input pixels with high weights identified by the attribution algorithm are marked by unit value in the mask matrix $M_{t}$, and the remaining pixels are assigned zero values. This mask matrix is then multiplied by the perturbation $\Delta_{t}$ to generate the adversarial $Patch_{t}$ (in Eq. 2). This step is useful as it can filter out non-important pixels in the inputs to reduce the number of perturbed pixels to improve the effectiveness of optimization-based attacks and reduce computation costs of runtime attacks.
@@
Finally, we develop a new initialization algorithm to shift the patch position and adjust its size when the detected BBox changes (Eq. 9-11). We shift the attack vector toward the new position of the detected BBox of the lead vehicle with a magnitude of ($x_{t}-x_{t-1}$, $y_{t}-y_{t-1}$), where ($x_{t-1}$, $y_{t-1}$) and ($x_{t}$, $y_{t}$) are the centers of BBox at previous and current control cycles, respectively (Eq. 9). We then expand the adversarial patch attack vector ($Patch$) to the dimension that matches the size of the newly detected BBox of the lead vehicle. Instead of reinitializing the whole attack vector matrix with random or zero values, which will reset the whole optimization process, we keep the previous patch values and intermediate variables and only initialize the newly expended units (Eq. 10-11). An example is shown in Fig. 5.
@@
IV Safety Intervention Simulation

### v4

@@ 3.1. Attack Model
A possible way to achieve this is to implant malware by compromising the over-the-air (OTA) update mechanisms (Cybersecurity, ; wen2020plug, ; mocnik_vehicular_nodate, ; elkhail2021vehicle, ) or gaining one-time remote access to the ADAS software through scanning the network, accessing stolen credentials and exploiting the vulnerabilities in SSH protocol (vulssh, ), browsers, access control (nie2017free, ), wireless communications (nie2017free, ; luo2022risks, ; hackjeep2015, ), third-party components connected to in-vehicular network (koscher2010experimental, ), or some remote service/backdoor offered by the manufacturer (e.g., Comma Connect for OpenPilot (comma-connect, ) or Bluelink for Hyundai). For example, a publicly-available tool developed for OpenPilot enables an attacker on the same network as a target device to install a malicious code (Workbench, ). With such remote access, the attacker can also change the OTA settings (e.g., remote URL) to prevent potential patches from being effective. This assumption about the attack surface for deploying malware is also supported by previous works (elkhail2021vehicle, ; eiza2017driving, ), and could have a large impact as it can be generalized to any vehicle with similar OTA and DNN mechanisms and target a large fleet of vehicles at the same time.
Table 2. Threat models: attacker strength, capability, and impact.
Threat Model Attacker Strength Access to ADAS Software Vehicular Networks Computation Location Impact Examples
Malware Strong1 $\checkmark$ r/w within ADAS
@@
(elkhail2021vehicle, ; eiza2017driving, )
Wireless Medium2 r/w
Local Device,
Remote Server
@@
(lagraa2019real, )(hackcamera, )(jha2020ml, )
(nie2017free, )(miller2015remote, )
Physical Weak3 r Remote Server
@@
(hoory2020dynamic, )(chahe2023dynamic, )
(lovisotto2021slap, )(man2023person, )(ma2023wip, )
@@
    Only perception attacks possible;
    r/w represents read (r) and write (w) access to vehicular networks.
Another way to compromise live camera data is to connect to a wireless communication device, either a third-party component or one implemented by an attacker, connected to the vehicular network, such as ROS communication channels (lagraa2019real, ), CAN Bus (nie2017free, ; openpilot_hardware, ; miller2015remote, ; wen2020plug, ) or Ethernet channel (jha2020ml, ; hackcamera, )), to read and send image data at runtime. The attacker computes the attack value on a local wireless device or a remote server.
Further, physical attack methods are also viable, such as by displaying the patch on a monitor attached on the rear side of a leading adversarial vehicle (hoory2020dynamic, ; chahe2023dynamic, ) or projecting the patch into the rear of the LV using a projector (lovisotto2021slap, ; man2023person, ; ma2023wip, ).
Table 2 summarizes various methods for runtime reading and modifying of live camera frames, given different attacker strengths and capabilities. In this paper, we mainly focus on the runtime and optimized modification of live camera frames to enhance attack success rate and stealthiness, regardless of the threat model and how the attacker obtained access. In our experiments, we implement the attack through malicious OTA update to OpenPilot (Section 3.3).
Attacker Constraints. We restrict the scope of attack capabilities in the paper to reading and modifying live camera data, ensuring uniformity across all presented threat models. Although some attack models could potentially enable more aggressive attacks, such as directly altering ACC controller outputs ( 3 in Fig. 1) via malware or wireless method or changing the DNN output ( 2) with malware method, these may be easier or earlier identified by safety mechanisms (e.g., AEBS) or human driver (see Section 5.4).
@@ 3.2. Attack Challenges
C1. Optimal timing of attacks at runtime to cause safety hazards. Prior attacks on ADAS that rely on random strategies to determine the attack timing (start time and duration) have proven ineffective in achieving a high attack success rate (zhou2022strategic, ; rubaiyat2018experimental, ; jha2019ml, ) as they waste computational resources by trying random attack parameters that lead to no safety hazards. For instance, initiating an attack on an Ego vehicle to induce sudden acceleration does not cause safety hazards when no LV is detected. Recent works have focused on using machine learning to explore the fault/attack parameter space (moradi2020exploring, ) and improve the attack effectivenes (jha2019ml, ; jha2020ml, ), but they still require substantial amounts of data from random attack experiments for model training. Finding the optimal triggering time and duration is crucial for effective attacks, yet challenging due to the vast parameter space that needs exploration.
C2. Generating attack value at runtime to adapt to dynamic changes in the driving environment. Attacking DNN-based ACC systems on a moving vehicle faces challenges due to continuous variations in the driving environment, such as object position and size captured by the Ego vehicle’s camera. Existing attack algorithms (eykholt2018robust, ; guo2023adversarial, ) are inadequate as they plan perturbations offline, assuming fixed sizes and locations for attack vectors. A new algorithm is needed to dynamically adapt the attack vector’s value (e.g., position, dimension, and amount of perturbation) to match the LV’s dynamics. These changes disrupt the original attack vector generation process, requiring a unique approach to address inconsistencies and non-differentiability in the objective function. In addition, the attack value should be designed in a stealthy way to avoid detection by the human driver or safety mechanisms.
C3. Incorporating real-time constraints into the attack optimization process. Previous attacks on DNN models assume predetermined target images (eykholt2018robust, ) or a known set (sato2021dirty, ; guo2023adversarial, ) with unlimited computation resources, allowing iterative optimization until an optimal attack vector is generated. However, attacking ACC systems in real-time presents challenges as the camera continuously provides frames without prior knowledge. An attack vector must be generated in real-time before the next frame or control action execution. The real-time control cycle and camera update frequency limit the speed of generating the adversarial attack vector and the frequency of assessing the perturbation’s impact on DNN predictions. These tight constraints in typical ACC systems (e.g., frame rate of 20Hz and control cycle of 10ms in (commaai-openpilot, )) significantly impact the effectiveness of optimization-based attack strategies.
3.3. Attack Design
@@
To tackle challenge C1, rather than randomly or exhaustively exploring the attack parameter space, we systematically characterize specific values within the parameter space (e.g., attack start time and duration). This targeted approach aims to identify optimal system contexts (or critical times) for activating the attacks to not waste time and resources on non-hazardous scenarios.
@@
To address the challenges C2-C3, we design a novel optimization approach and an adaptive algorithm to dynamically determine optimal pixel values for an adversarial patch at runtime, aiming to maximize the error in DNN-based LV predictions and to accommodate dynamic changes in the driving environment. To optimize attack effectiveness and computational efficiency, our method focuses on the small area of the bounding box around the target vehicle and employs a primary attribution algorithm (sundararajan2017axiomatic, ) to identify and manipulate the most crucial pixels (Section 3.3.2). Additionally, our adaptive algorithm retains optimization results of perturbation size, position, and value from the previous perception cycle instead of restarting the optimization process (contrary to other iterative optimization methods (croce2020reliable, )) to satisfy the tight real-time constraints of the perception system (50ms).
Figure 2. Attack design: Offline preparation, Runtime execution.
@@
The attacker initiates the attack process by conducting an offline analysis on the target DNN-based ACC system. The analysis includes examining operational data and open-source code to identify the target software files and functions and/or the DNN input and output fields to be monitored and infected. Then the attack steps outlined in the orange box in Fig. 2, are implemented and executed either remotely or locally through unauthorized access to the target ACC under one of the threat models described in Table 2. In Fig 2, we present an attack implementation example via malware method that will replace the target functions or system libraries. The attacker installs the malware on the target ACC system through one-time access to the victim vehicle control system, achieved by exploiting remote access vulnerabilities or compromising OTA updates (Section 3.1). Appendix B provides an illustrative example of how we install the malware on an OpenPilot system. At runtime, the malicious code intercepts sensor data before reaching the DNN model and infers the current system context for activating attacks. When the detected system context aligns with a critical system context (Section 3.3.1), an optimized adversarial patch is generated (Section 3.3.2) and added to the image data and sent to DNN model.
3.3.1. Context-Aware Attack Activation
@@
To find the most critical times for activation of the attack (C1), we adopt a control-theoretic hazard analysis method (leveson2013stpa, ) to identify the most critical system contexts under which specific control actions are unsafe and, if issued by the ACC control software, could lead to hazards. This approach mainly relies on domain knowledge about system safety requirements and, contrary to an ML-based approach, does not require large amounts of training data or computation resources.
@@
In our hazard analysis, we define the accident as the adverse event of forward collision (measured by a zero or negative relative distance between the Ego vehicle and the LV or a front object). This can happen as a result of the Ego vehicle transitioning into a hazardous state that violates the safe following distance with LV.
@@
To determine the critical system contexts, we assess all the combinations of system states and ACC control actions (e.g., acceleration, deceleration) to identify the specific combinations that are most likely to lead to hazards. Table 3 shows part of the safety context table for ACC with a focus on the acceleration commands. Here, the critical system context is described as when the Headway Time (HWT, the time the Ego vehicle takes to drive the relative distance ($RD$) from the LV with the current speed) is less than a safety limit HWTsafe (e.g., 2-3s), and the Ego vehicle is faster than the LV ($V_{Ego}>V_{Lead}$). Under such system context, an acceleration command induced by the camera perception attacks or other reasons will most likely lead to a forward collision hazard. This high-level specification of critical system context can be done by an attacker based on the knowledge of the typical functionality of an ACC system and be applied to any ACC system with the same functional specification.
Table 3. Partial safety context table for an ACC system.
@@ 4 RS>0 No
3.3.2. Adaptive Adversarial Patch Generation
@@
(1) min $\sum_{d\in{RD}_{t}}{-\triangledown g(d,\theta)}+\lambda||\Delta_{t}||_{p}$
(2) s.t. $Patch_{t}=\Delta_{t}*M_{t}$
(3) $Patch_{t}\in[\mu-\sigma,\mu+\sigma]$
(4) $Area(Patch_{t})\subset BBox(LV)_{t}$
(5) $X^{adv}_{t}=X_{t}+Patch_{t}$
(6) $[{RD},RS]_{t}=LVD_{\theta}(X^{adv}_{t-1})$
(7) $u_{t}=ACC(s_{t},[{RD},{RS}]_{t})$
(8) $s_{t+1}=CarModel(s_{t},u_{t})$
@@
Figure 3. Optimization-based adversarial patch generation.
@@
In Eq. 1, $g(d)$ is an approximate polynomial function of $d$ that fits the trend of the trajectory of the relative distance ${RD}_{t}$, predicted by the DNN model with weight parameters $\theta$. "-" is a negative sign that converts our goal of maximizing the relative distance to minimizing the proposed objective function. For example, when the gradient of the relative distance trajectory, $g(d)$, is negative, minimizing "$-\triangledown g(d)$" will slow down the decrease of $g(d)$ and is equivalent to maximizing the relative distance. We adopt the gradient of $g(d)$ in the objective function instead of using ${RD}_{t}$ itself in order to avoid sharp changes in the predicted relative distance value, which might be easily detected by some anomaly detection mechanisms. We assume the attacker has access to the DNN predictions (e.g., $RD$) by monitoring the ADAS communication network (e.g., ROS) or by running a replicated DNN model on a remote server or wireless communication device (see Table 2).
@@
In Eq. 5, the perturbation is added to the original image input $X_{t}\in\mathbb{R}^{H\times W\times C}$ in the form of an adversarial patch $Patch_{t}\in\mathbb{R}^{H\times W\times C}$, represented as a matrix of pixels with height H, width W, and C color channels. $\lambda$ is the weight parameter of the p-norm regularization term, designed to minimize the perturbation value of the patch for stealthiness. We limit the perturbation value within the Kalman filter noise parameters ($\mu$,$\sigma$) (Eq. 3), which ensures the perturbation is not corrected by the sensor fusion. We also constrain the adversarial patch inside the BBox of the LV (Eq. 4) to enhance attack effectiveness, minimize the perturbation area for stealthiness, and reduce computational cost.
@@
Figure 4. Examples of the shift and adjustment process in the patch generation. Inset figures are the zoomed-in views of the front vehicle with an adversarial patch added around the license plate area.
Figure 5. ACC under attack.
@@
Primary Attribution Detection and Patch Update. As mentioned in Section 3.2, a major challenge (C3) in the design of runtime attacks is the changes in the size and location of the LV in the perceived image frames. To address this challenge, the attacker needs to update the generated patch dynamically according to the approximate DNN outputs. In this paper, we adopt an object detection method (redmon2016you, ) to detect and track the real-time location and size of the LV and concentrate the attack perturbation within the detected BBox of the LV. In production ACC with object detection features (Apollo, ), given proper access, the attacker can skip this step and use the stock prediction results.
@@
After getting the BBox, we utilize a primary attribution algorithm (sundararajan2017axiomatic, ) to quantify the relationship between input features and output predictions. Through this exploration, we try to identify the important pixels within the BBox of the LV that contribute the most to the predictions of $RD_{t}$. The input pixels with high weights identified by the attribution algorithm are marked by unit value in the mask matrix $M_{t}$, and the remaining pixels are assigned zero values. This mask matrix is then multiplied by the perturbation $\Delta_{t}$ to generate the adversarial $Patch_{t}$ (in Eq. 2). This step is useful as it can filter out non-important pixels in the inputs to reduce the number of perturbed pixels to improve the effectiveness of optimization-based attacks and reduce computation costs of runtime attacks.
@@
Finally, we develop a new initialization algorithm to shift the patch position and adjust its size when the detected BBox changes (Eq. 9-11). We shift the attack vector toward the new position of the detected BBox of the LV with a magnitude of ($x_{t}-x_{t-1}$, $y_{t}-y_{t-1}$), where ($x_{t-1}$, $y_{t-1}$) and ($x_{t}$, $y_{t}$) are the centers of BBox at previous and current control cycles, respectively (Eq. 9). We then expand the adversarial patch attack vector ($Patch$) to the dimension that matches the size of the newly detected BBox of the LV. Instead of reinitializing the whole attack vector matrix with random or zero values, which will reset the whole optimization process, we keep the previous patch values and intermediate variables and only initialize newly expended units (Eq. 10-11). Fig. 5 shows an example.
@@
4. Safety Intervention Simulation